OpenStudy (anonymous):

Find the first five terms of the given sequence Please Help (Sequence in comments)

OpenStudy (anonymous):

\[a _{1}=9,a _{n+1}=\frac{ 2a _{n}-2 }{ 2 }\]

OpenStudy (anonymous):

a.9, 7, 5, 3, 1 b.9, 16, 14, 12, 10 c.9, 8, 7, 6, 5 d.9, 1, 2, 3, 4

OpenStudy (anonymous):

:3

OpenStudy (ranga):

First simplify \[a _{n+1}=\frac{ 2a _{n}-2 }{ 2 }\]

OpenStudy (ranga):

\[a _{n+1}=\frac{ 2a _{n}-2 }{ 2 } = a_n - 1\]

OpenStudy (ranga):

So the next term is just the previous term minus 1.

OpenStudy (anonymous):

\[a _{2}=\frac{ 2a _{1}-2 }{2 }\]?

OpenStudy (anonymous):

oh so that would make it 9,8,7,6,5?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

Thanks!

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

(✌゚∀゚)☞